# CP-Guard: A Unified, Probability-Agnostic, and Adaptive Framework for Malicious Agent Detection and Defense in Multi-Agent Embodied Perception Systems

### v1

@@ Abstract
Collaborative Perception (CP) has been shown to be a promising technique for multi-agent autonomous driving and multi-agent robotic systems, where multiple agents share their perception information to enhance the overall perception performance and expand the perception range. However, in CP, an ego agent needs to receive messages from its collaborators, which makes it vulnerable to attacks from malicious agents. To address this critical issue, we propose a unified, probability-agnostic, and adaptive framework, namely, CP-Guard, which is a tailored defense mechanism for CP deployed by each agent to accurately detect and eliminate malicious agents in its collaboration network. Our key idea is to enable CP to reach a consensus rather than a conflict against an ego agent’s perception results. Based on this idea, we first develop a probability-agnostic sample consensus (PASAC) method to effectively sample a subset of the collaborators and verify the consensus without prior probabilities of malicious agents. Furthermore, we define collaborative consistency loss (CCLoss) for object detection task and bird’s eye view (BEV) segmentation task to capture the discrepancy between an ego agent and its collaborators, which is used as a verification criterion for consensus. In addition, we propose online adaptive threshold via dual sliding windows to dynamically adjust the threshold for consensus verification and ensure the reliability of the systems in dynamic environments. Finally, we conduct extensive experiments and demonstrate the effectiveness of our framework. Code will be released at https://github.com/CP-Security/CP-Guard.
@@ 1 Introduction
Figure 1: Illustration of the threats of malicious agent in collaborative perception and our defense framework, CP-Guard. When there is no defense, malicious agents could easily send intricately crafted adversarial messages to an ego agent, consequently misleading the CP system to yield false perception outputs. To counter this vulnerability, we propose CP-Guard, a tailored defense mechanism for CP to effectively detect and neutralize malicious agents, thereby ensuring robust perception outcomes.
@@
In order to address this issue, several works have explored this problem. For example, Li et al. [5] proposed robust collaborative sampling consensus (ROBOSAC) to randomly sample a subset of the collaborators and verify the consensus, but it requires the prior probabilities of malicious agents, which are usually unknown in practice. In addition, Zhao et al. [6] and Zhang et al. [7] also developed defense methods against malicious agents, while these methods need to check the collaborators one by one, which is inefficient and computation-intensive. Moreover, other works [4, 8, 9] used adversarial training to enhance the robustness of the model. However, adversarial training introduces additional overhead during training and lacks generalization for unseen attacks. Additionally, it may result in a reduction in accuracy and is non-trivial to achieve computationally efficient and generalizable adversarial defense in CP.
In order to overcome the aforementioned limitations, we design a novel defense mechanism, CP-Guard. It can be deployed by each agent to accurately detect and eliminate malicious agents in its local collaboration network. The key idea is to enable CP to achieve a consensus rather than a conflict against an ego agent’s perception results. Following this idea, we first design a probability-agnostic sample consensus (PASAC) method to effectively sample a subset of the collaborators and verify the consensus without prior probabilities of malicious agents. In addition, the consensus is verified by our carefully designed collaborative consistency loss (CCLoss), which is used to calculate the discrepancy between an ego agent and its collaborators. If a collaborator’s collaborative consistency loss exceeds a certain threshold, the collaborator is considered to be a benign agent, otherwise, it is considered to be a malicious agent. In order to ensure a low false positive rate and the reliability of the system, we propose an online adaptive threshold to dynamically adjust the threshold for consensus verification. The main contributions of this paper are summarized as follows.
  • We analyze the vulnerabilities of CP against malicious agents and develop a novel framework for robust collaborative BEV perception, CP-Guard, which can defend against attacks and eliminate malicious agents from a local collaboration network.
@@
  • Finally, we conduct extensive experiments on collaborative object detection and collaborative BEV segmentation tasks and demonstrate the effectiveness of our CP-Guard.
The remainder of this paper is organized as follows. In Section 2, we review the related works on collaborative perception, adversarial perception, and defensive perception. In Section 3, we provide the formulation and pipeline of collaborative perception, and then we discuss the adversarial threat model. In Section 4, we propose our unified, probability-agnostic, and adaptive framework, CP-Guard, which consists of a probability-agnostic sample consensus (PASAC) method and two collaborative consistency losses (CCLoss) for the object detection task and BEV segmentation task. In addition, we also propose an online adaptive threshold scheme to dynamically adjust the threshold for consensus verification, which can ensure the reliability of the system in dynamic environments. In Section 5, we conduct extensive experiments to evaluate the performance of our framework. Finally, we discuss the limitations and future work in Section 6 and conclude this paper in Section 7.
@@ 2.3 Defensive Perception
To fortify intermediate-level collaborative perception systems against adversarial attacks, Li et al. [5] proposed the Robust Collaborative Sampling Consensus (ROBOSAC) method. This approach entails a random selection of a subset of collaborators for consensus verification. Despite its potential, the efficacy of ROBOSAC hinges on the availability of prior probabilities of malicious intent among agents, which are often unknown in real-world scenarios. Moreover, Zhao et al. [6] and Zhang et al. [7] have formulated defensive strategies identifying malicious agents. These techniques, however, involve scrutinizing each collaborator individually, which are both computationally intensive and inefficient. Adversarial training has also been explored as a mechanism to bolster system robustness, as demonstrated in the studies by Tu et al. [4], Raghunathan et al. [8], and Zhang et al. [9]. While this approach enhances system robustness to address security, it substantially increases the computational load during training and may not effectively generalize to novel, unseen attacks [28, 29, 30, 31, 32]. Additionally, adversarial training often results in diminished model accuracy and poses significant challenges in developing a computationally efficient and scalable adversarial defense that can be broadly applied across collaborative perception platforms. In contrast, our approach introduces a methodology that can be autonomously implemented by each agent to accurately detect and neutralize malicious entities within the local collaborative network, aiming to enhance both the efficiency and effectiveness of defense mechanisms in collaborative perception systems.
@@ 3.2 Adversarial Threat Model in Collaborative Perception
4 CP-Guard Framework
In this section, we present our CP-Guard in detail. It consists of three main components: (1) Probability-Agnostic Sample Consensus (PASAC), (2) Collaborative Consistency Loss Verification (CCLoss), and (3) Online Adaptive Threshold. PASAC is designed to effectively sample a subset of collaborators for consensus verification without relying on prior probabilities of malicious intent. CCLoss is proposed to verify the consensus between the ego agent and the collaborative agents. Online Adaptive Threshold is leveraged to dynamically adjust the threshold for consensus verification. These three components work collaboratively to detect and neutralize malicious agents in the local collaboration network. We elaborate on these three components in the following sections.
@@ 4.3 Online Adaptive Threshold via Dual Sliding Windows
Algorithm 3 Dual‑Window Online Threshold Update
2:if $\hat{y}_{t}=\mathtt{BENIGN}$ then
3:     push_back$(\mathcal{W}^{P},Z_{t})$
4:else
5:     push_back$(\mathcal{W}^{N},Z_{t})$
6:end if
7:truncate both windows to length $W$
8:if $|\mathcal{W}^{P}|\!\geq\!W_{\min}$ and $|\mathcal{W}^{N}|\!\geq\!W_{\min}$ then
9:     compute $\hat{q}_{1-\alpha}^{P},\hat{q}_{\beta}^{N}$
12:end if
13:return $\varepsilon$
@@ 5 Experiments
CP-Guard (against FGSM attack) 52.76 46.35 46.67 68.32 28.98 20.51 40.15 39.30
CP-Guard (against C&W attack) 49.22 44.08 44.76 65.58 30.12 20.83 39.10 37.95
CP-Guard (against PGD attack) 52.84 46.41 46.73 68.41 29.01 20.48 40.16 39.34
@@ 5.2 Evaluation and Analysis
5.2.1 Evaluation of CP-Guard.
We evaluate the efficacy of our CP-Guard scheme against a variety of adversarial attacks. The outcomes of these evaluations are detailed in Table I and Table II. As shown in Table I, in scenarios where the CP system lacks defensive mechanism, the mIoU across all three attack modalities significantly falls below the established lower bound, registering at 37.09. This substantial degradation in performance underscores the effectiveness of the adversarial attacks implemented. Conversely, our CP-Guard framework can effectively defend against these attacks and achieve an mIoU that closely approaches the upper bound of 40.45.
Additionally, we evaluate our approach on the object detection task. Table II presents the Average Precision (AP) at thresholds 0.5 and 0.7 for different defense strategies. Our observations indicate that CP-Guard consistently achieves the highest AP@0.5 and AP@0.7 across all attack scenarios, demonstrating its superior performance. Specifically, when defending against the PGD attack, CP-Guard achieves an AP@0.5 of 80.4 and an AP@0.7 of 78.3, which is higher than ROBOSAC (77.9 and 75.6) by 2.5 point and 2.7 point, respectively. For defending against the C&W attack, CP-Guard performs better, achieving an AP@0.5 of 80.2 and an AP@0.7 of 77.6, which is higher than ROBOSAC (74.5 and 71.1) by 5.7 point and 6.5 point, respectively.
These quantitative results confirm that CP-Guard not only defends BEV segmentation (mIoU) but also generalizes effectively to object detection, restoring performance close to the all-benign scenario and substantially outperforming both adversarially trained baselines and prior-art methods such as ROBOSAC.
@@
CP-Guard (against PGD attack) 80.4 78.3
@@
CP-Guard (against C&W attack) 80.2 77.6
@@
Figure 5: Visualization of no defense and defensive CP-Guard results on V2X-Sim datasets.
@@ 5.2.5 Qualitative Evaluation
As depicted in Fig. 5, we present the visualization results on the V2X-Sim dataset. Without CP-Guard, attackers can significantly disrupt collaborative perception, leading to a marked degradation in the performance of BEV segmentation tasks. However, our introduced CP-Guard framework can intelligently identify benign collaborators and eliminate malicious collaborators, thereby facilitating robust CP.
@@ 6 Limitations and Future Work
Despite the promising results of CP-Guard in defending against malicious agents in collaborative perception, the computational complexity of PASAC, though more efficient than ROBOSAC and linear sampling which checking every agent individually, still introduces $\mathcal{O}(N\log N)$ overhead in the worst case. This may challenge real-time performance requirements in dense traffic scenarios with numerous collaborative agents. In the future, we will explore more lightweight pre-screening methods to further reduce computational requirements of PASAC in dense collaborative networks, and expand the defensive capabilities against more sophisticated adversarial attacks by incorporating temporal consistency checks that leverage sequential data to identify anomalous behavior patterns.
@@ 7 Conclusion
In this paper, we have designed a novel defense framework for CP named CP-Guard, which consists of three parts. The first is PASAC which can effectively sample the collaborators without the prior probabilities of malicious agents. The second is collaborative consistency loss verification which calculates the discrepancy between the ego agent and the collaborators, which is used as a verification criterion for consensus. The third is the online adaptive threshold which can adaptively adjust the threshold to ensure the stability of the system in dynamic environments. Extensive experiments show that our CP-Guard can defend against different types of attacks and can adaptively adjust the trade-off between performance and computational overhead.

### v2

# CP-uniGuard: A Unified, Probability-Agnostic, and Adaptive Framework for Malicious Agent Detection and Defense in Multi-Agent Embodied Perception Systems

@@ Abstract
Collaborative Perception (CP) has been shown to be a promising technique for multi-agent autonomous driving and multi-agent robotic systems, where multiple agents share their perception information to enhance the overall perception performance and expand the perception range. However, in CP, an ego agent needs to receive messages from its collaborators, which makes it vulnerable to attacks from malicious agents. To address this critical issue, we propose a unified, probability-agnostic, and adaptive framework, namely, CP-uniGuard, which is a tailored defense mechanism for CP deployed by each agent to accurately detect and eliminate malicious agents in its collaboration network. Our key idea is to enable CP to reach a consensus rather than a conflict against an ego agent’s perception results. Based on this idea, we first develop a probability-agnostic sample consensus (PASAC) method to effectively sample a subset of the collaborators and verify the consensus without prior probabilities of malicious agents. Furthermore, we define collaborative consistency loss (CCLoss) for object detection task and bird’s eye view (BEV) segmentation task to capture the discrepancy between an ego agent and its collaborators, which is used as a verification criterion for consensus. In addition, we propose online adaptive threshold via dual sliding windows to dynamically adjust the threshold for consensus verification and ensure the reliability of the systems in dynamic environments. Finally, we conduct extensive experiments and demonstrate the effectiveness of our framework. Code will be released at https://github.com/CP-Security/CP-uniGuard.
@@ 1 Introduction
Figure 1: Illustration of the threats of malicious agent in collaborative perception and our defense framework, CP-uniGuard. When there is no defense, malicious agents could easily send intricately crafted adversarial messages to an ego agent, consequently misleading the CP system to yield false perception outputs. To counter this vulnerability, we propose CP-uniGuard, a tailored defense mechanism for CP to effectively detect and neutralize malicious agents, thereby ensuring robust perception outcomes.
@@
In order to address this issue, several works have explored this problem. For example, Li et al. [5] proposed robust collaborative sampling consensus (ROBOSAC) to randomly sample a subset of the collaborators and verify the consensus, but it requires the prior probabilities of malicious agents, which are usually unknown in practice. In addition, Zhao et al. [6] and Zhang et al. [7] also developed defense methods against malicious agents, while these methods need to check the collaborators one by one, which is inefficient and computation-intensive. Moreover, Hu et al. [8] designed CP-Guard+, which trains an anomaly detector to identify malicious agents at the feature level. However, it requires collecting feature maps from both benign and malicious agents as training data, which is hard to achieve in real-world applications. Additionally, training an anomaly detector demands extra computational resources and lacks generalization for unseen attacks. For example, malicious agents may develop new attacks not covered by the training data to bypass the anomaly detector.
In order to overcome the aforementioned limitations, we design a novel defense mechanism, CP-uniGuard. It can be deployed by each agent to accurately detect and eliminate malicious agents in its local collaboration network. The key idea is to enable CP to achieve a consensus rather than a conflict against an ego agent’s perception results. Following this idea, we first design a probability-agnostic sample consensus (PASAC) method to effectively sample a subset of the collaborators and verify the consensus without prior probabilities of malicious agents. In addition, the consensus is verified by our carefully designed collaborative consistency loss (CCLoss), which is used to calculate the discrepancy between an ego agent and its collaborators. If a collaborator’s collaborative consistency loss exceeds a certain threshold, the collaborator is considered to be a benign agent, otherwise, it is considered to be a malicious agent. In order to ensure a low false positive rate and the reliability of the system, we propose an online adaptive threshold to dynamically adjust the threshold for consensus verification. The main contributions of this paper are summarized as follows.
  • We analyze the vulnerabilities of CP against malicious agents and develop a novel framework for robust collaborative BEV perception, CP-uniGuard, which can defend against attacks and eliminate malicious agents from a local collaboration network.
@@
  • Finally, we conduct extensive experiments on collaborative object detection and collaborative BEV segmentation tasks and demonstrate the effectiveness of our CP-uniGuard.
The remainder of this paper is organized as follows. In Section 2, we review the related works on collaborative perception, adversarial perception, and defensive perception. In Section 3, we provide the formulation and pipeline of collaborative perception, and then we discuss the adversarial threat model. In Section 4, we propose our unified, probability-agnostic, and adaptive framework, CP-uniGuard, which consists of a probability-agnostic sample consensus (PASAC) method and two collaborative consistency losses (CCLoss) for the object detection task and BEV segmentation task. In addition, we also propose an online adaptive threshold scheme to dynamically adjust the threshold for consensus verification, which can ensure the reliability of the system in dynamic environments. In Section 5, we conduct extensive experiments to evaluate the performance of our framework. Finally, we discuss the limitations and future work in Section 6 and conclude this paper in Section 7.
@@ 2.3 Defensive Perception
To fortify intermediate-level collaborative perception systems against adversarial attacks, Li et al. [5] proposed the Robust Collaborative Sampling Consensus (ROBOSAC) method. This approach entails a random selection of a subset of collaborators for consensus verification. Despite its potential, the efficacy of ROBOSAC hinges on the availability of prior probabilities of malicious intent among agents, which are often unknown in real-world scenarios. Moreover, Zhao et al. [6] and Zhang et al. [7] have formulated defensive strategies identifying malicious agents. These techniques, however, involve scrutinizing each collaborator individually, which are both computationally intensive and inefficient. Adversarial training has also been explored as a mechanism to bolster system robustness, as demonstrated in the studies by Tu et al. [4], Raghunathan et al. [27], and Zhang et al. [28]. While this approach enhances system robustness to address security, it substantially increases the computational load during training and may not effectively generalize to novel, unseen attacks [29, 30, 31]. Additionally, adversarial training often results in diminished model accuracy and poses significant challenges in developing a computationally efficient and scalable adversarial defense that can be broadly applied across collaborative perception platforms. In contrast, our approach introduces a methodology that can be autonomously implemented by each agent to accurately detect and neutralize malicious entities within the local collaborative network, aiming to enhance both the efficiency and effectiveness of defense mechanisms in collaborative perception systems.
@@ 3.2 Adversarial Threat Model in Collaborative Perception
4 CP-uniGuard Framework
In this section, we present our CP-uniGuard in detail. It consists of three main components: (1) Probability-Agnostic Sample Consensus (PASAC), (2) Collaborative Consistency Loss Verification (CCLoss), and (3) Online Adaptive Threshold. PASAC is designed to effectively sample a subset of collaborators for consensus verification without relying on prior probabilities of malicious intent. CCLoss is proposed to verify the consensus between the ego agent and the collaborative agents. Online Adaptive Threshold is leveraged to dynamically adjust the threshold for consensus verification. These three components work collaboratively to detect and neutralize malicious agents in the local collaboration network. We elaborate on these three components in the following sections.
@@ 4.3 Online Adaptive Threshold via Dual Sliding Windows
Algorithm 3 Dual‑Window Online Threshold Update
2:if $\hat{y}_{t}=\mathtt{BENIGN}$ then
3:     push_back$(\mathcal{W}^{P},Z_{t})$
4:else
5:     push_back$(\mathcal{W}^{N},Z_{t})$
6:end if
7:truncate both windows to length $W$
8:if $|\mathcal{W}^{P}|\!\geq\!W_{\min}$ and $|\mathcal{W}^{N}|\!\geq\!W_{\min}$ then
9:     compute $\hat{q}_{1-\alpha}^{P},\hat{q}_{\beta}^{N}$
12:end if
13:return $\varepsilon$
@@ 5 Experiments
CP-uniGuard (against FGSM attack) 52.76 46.35 46.67 68.32 28.98 20.51 40.15 39.30
CP-uniGuard (against C&W attack) 49.22 44.08 44.76 65.58 30.12 20.83 39.10 37.95
CP-uniGuard (against PGD attack) 52.84 46.41 46.73 68.41 29.01 20.48 40.16 39.34
@@ 5.2 Evaluation and Analysis
5.2.1 Evaluation of CP-uniGuard.
We evaluate the efficacy of our CP-uniGuard scheme against a variety of adversarial attacks. The outcomes of these evaluations are detailed in Table I and Table II. As shown in Table I, in scenarios where the CP system lacks defensive mechanism, the mIoU across all three attack modalities significantly falls below the established lower bound, registering at 37.09. This substantial degradation in performance underscores the effectiveness of the adversarial attacks implemented. Conversely, our CP-uniGuard framework can effectively defend against these attacks and achieve an mIoU that closely approaches the upper bound of 40.45.
Additionally, we evaluate our approach on the object detection task. Table II presents the Average Precision (AP) at thresholds 0.5 and 0.7 for different defense strategies. Our observations indicate that CP-uniGuard consistently achieves the highest AP@0.5 and AP@0.7 across all attack scenarios, demonstrating its superior performance. Specifically, when defending against the PGD attack, CP-uniGuard achieves an AP@0.5 of 80.4 and an AP@0.7 of 78.3, which is higher than ROBOSAC (77.9 and 75.6) by 2.5 point and 2.7 point, respectively. For defending against the C&W attack, CP-uniGuard performs better, achieving an AP@0.5 of 80.2 and an AP@0.7 of 77.6, which is higher than ROBOSAC (74.5 and 71.1) by 5.7 point and 6.5 point, respectively.
These quantitative results confirm that CP-uniGuard not only defends BEV segmentation (mIoU) but also generalizes effectively to object detection, restoring performance close to the all-benign scenario and substantially outperforming both adversarially trained baselines and prior-art methods such as ROBOSAC.
@@
CP-uniGuard (against PGD attack) 80.4 78.3
@@
CP-uniGuard (against C&W attack) 80.2 77.6
@@
Figure 5: Visualization of no defense and defensive CP-uniGuard results on V2X-Sim datasets.
@@
TABLE IV: Generalization Evaluation. We evaluate and compare the generalization capability of CP-uniGuard on two attacks, PGD and C&W, respectively.
Method AP@0.5 AP@0.7
CP-Guard+ (against PGD attack) 71.7 68.6
CP-uniGuard (against PGD attack) 80.4 (+8.7) 78.3 (+9.7)
CP-Guard+ (against C&W attack) 73.9 71.2
CP-uniGuard (against C&W attack) 80.2 (+6.3) 77.6 (+6.4)
5.2.5 Generalization Evaluation
We further evaluate the generalization capability of CP-uniGuard across different attack types, comparing it with feature-level defense methods such as CP-Guard+ [8]. For CP-Guard+, we train the model on the CP-GuardBench dataset [8] while excluding specific attacks from the training set to assess its generalization performance. Specifically, we exclude the PGD attack and C&W attack during training and subsequently evaluate the model’s performance against these two attacks, respectively. Results are presented in Table IV, we can see that CP-uniGuard consistently achieves superior AP@0.5 and AP@0.7 scores compared to CP-Guard+ under both PGD and C&W attacks which are unseen. This performance gap stems from CP-uniGuard’s fundamental design: rather than relying on training, CP-uniGuard leverages consensus verification, which provides inherent generalization to unseen attacks without requiring prior exposure during training.
@@ 5.2.6 Qualitative Evaluation
As depicted in Fig. 5, we present the visualization results on the V2X-Sim dataset. Without CP-uniGuard, attackers can significantly disrupt collaborative perception, leading to a marked degradation in the performance of BEV segmentation tasks. However, our introduced CP-uniGuard framework can intelligently identify benign collaborators and eliminate malicious collaborators, thereby facilitating robust CP.
@@ 6 Limitations and Future Work
Despite the promising results of CP-uniGuard in defending against malicious agents in collaborative perception, the computational complexity of PASAC, though more efficient than ROBOSAC and linear sampling which checking every agent individually, still introduces $\mathcal{O}(N\log N)$ overhead in the worst case. This may challenge real-time performance requirements in dense traffic scenarios with numerous collaborative agents. In the future, we will explore more lightweight pre-screening methods to further reduce computational requirements of PASAC in dense collaborative networks, and expand the defensive capabilities against more sophisticated adversarial attacks by incorporating temporal consistency checks that leverage sequential data to identify anomalous behavior patterns.
@@ 7 Conclusion
In this paper, we have designed a novel defense framework for CP named CP-uniGuard, which consists of three parts. The first is PASAC which can effectively sample the collaborators without the prior probabilities of malicious agents. The second is collaborative consistency loss verification which calculates the discrepancy between the ego agent and the collaborators, which is used as a verification criterion for consensus. The third is the online adaptive threshold which can adaptively adjust the threshold to ensure the stability of the system in dynamic environments. Extensive experiments show that our CP-uniGuard can defend against different types of attacks and can adaptively adjust the trade-off between performance and computational overhead.